OpenStudy (charlotte123):

Choose the equation of the line passing through the point (3, -2) and parallel to y = 3x - 6. y = 3x + 7 y = 3x - 11 y = -3x + 11 y = -3x - 12

hartnn (hartnn):

can u find slope of = 3x - 6. ?

OpenStudy (charlotte123):

no I tried but I cant D:

hartnn (hartnn):

slope of y=mx+c is just m

OpenStudy (charlotte123):

so would the answer be d? :P

hartnn (hartnn):

if you compare y= = 3x - 6. with y=mx+c you get m=3 so , slope =3 and parallel lines have same slope, so your required line will be of the form y=3x+b

hartnn (hartnn):

now just put x=3, y=-2 in y= 3x+b and find 'b'

OpenStudy (charlotte123):

oh I see now :O

hartnn (hartnn):

can u find b ?

OpenStudy (charlotte123):

b = -11 :D

hartnn (hartnn):

\(\checkmark \) good :) so whch option is it ?

OpenStudy (charlotte123):

b :D

OpenStudy (charlotte123):

thanks so much for the help! :D

hartnn (hartnn):

welcome ^_^